# An unidentified Fermi source emitting radio bursts in the Galactic bulge

### v1

@@ Abstract
We report on the detection of radio bursts from the Galactic bulge using the real-time transient detection and localization system, realfast. The pulses were detected commensally on the Karl G. Jansky Very Large Array during a survey of unidentified Fermi -ray sources. The bursts were localized to subarcsecond precision using realfast fast-sampled imaging. Follow-up observations with the Green Bank Telescope detected additional bursts from the same source. The bursts do not exhibit periodicity in a search up to periods of 480s, assuming a duty cycle of 20%. The pulses are nearly 100% linearly polarized, show circular polarization up to 12%, have a steep radio spectral index of –2.7, and exhibit variable scattering on timescales of months. The arcsecond-level realfast localization links the source confidently with the Fermi -ray source and places it nearby (though not coincident with) an XMM-Newton X-ray source. Based on the source’s overall properties, we discuss various options for the nature of this object and propose that it could be a young pulsar, magnetar, or a binary pulsar system.
@@ 1 Introduction
Astrophysical transients are events whose duration ranges between milliseconds to years. They inform us about the most variable and energetic events in the Universe. Radio transients are typically described as “fast” if they last for one second (Law et al., 2018a). Some of the fast radio transients, including pulsars (Hewish et al., 1968), rotating radio transients (RRATs) (McLaughlin et al., 2006), magnetar bursts, and fast radio bursts (FRBs) (Lorimer et al., 2007), are millisecond-duration emissions that are energetic enough to invoke coherent emission mechanisms to describe their high brightness temperature ( K) and luminosities. Most of these are produced by highly magnetized neutron stars, even though the origin of FRBs is highly debated. Pulsars, RRATs, and magnetars have a rotational period, while FRBs do not. FRBs are also many orders of magnitude brighter and more energetic than Galactic pulsars, magnetars, or RRATs. Nevertheless, these transients are unique probes in understanding the baryonic content of the Milky Way interstellar medium (ISM) and the intergalactic medium (IGM), in the case of FRBs (Macquart et al., 2020), along their line-of-sight (LOS). Regular timing of millisecond pulsars (MSPs) has been used to detect stochastic gravitational wave background in the Universe (Hellings & Downs, 1983; Agazie et al., 2023). Multiple ways have been proposed to detect these radio transients, including all-sky blind searches. One such method is to look for radio counterparts in the known high-energy sources, as some of these sources, like pulsars and magnetars, are established high-energy emitters.
@@
Fermi is a space mission that studies the cosmos in the energy range 10 keV–300 GeV. The imaging telescope on Fermi called the Large Area Telescope (LAT) has a wide field of view (8000 cm) and an angular resolution of . Fermi-LAT (4FGL-DR4) reported about 7195 point sources in their latest data release (Ballet et al., 2023). About 2427 (one-third) of these Fermi sources don’t have counterparts in any other electromagnetic regime. The origin and nature of these unassociated sources remain a mystery. A large fraction of these are eventually expected to be associated with the largest associated source classes, active galactic nuclei or pulsars. Recently, there appears a third class of so-called soft Galactic unassociated sources that are found in high-density regions of the Galactic plane. For a more detailed discussion of the unassociated gamma-ray source population (see Abdollahi et al. (2022)). Fermi-LAT has detected at least 297 -ray pulsars, 70% of which are radio-loud, and MSPs and young pulsars dominate this sample (Smith et al., 2023). Therefore, searching for new pulsars in the Fermi sources remains a reasonable case.
@@
The wide field of view and precise localization capabilities make radio interferometers ideal for searching for radio transients. Radio imaging-based pulsar searches have led to the discovery of the first isolated MSPs (Erickson, 1980) and the first globular cluster pulsar (Hamilton et al., 1985). This technique is based on the assumption that pulsars are compact and steep spectrum radio sources. Such searches have successfully led to the discovery of many MSPs and normal pulsars (Frail et al., 2016; Bhakta et al., 2017).
@@
realfast is a real-time commensal transient search system deployed at the Karl G. Janksy Very Large Array (JVLA) (Law et al., 2015, 2018b). It makes real-time interferometric images on visibility data sampled at 10 ms duration. This has the advantage of simultaneously detecting and localizing the transient to sub-arcsecond precision. realfast has so far played a crucial role in the localization and host galaxy determination of many FRBs, including the first repeating FRB 20121102A (Chatterjee et al., 2017), FRB 20180916B (Aggarwal et al., 2020), FRB 20180301A (Bhandari et al., 2022), FRB 20201124A (Ravi et al., 2022), FRB 20200120E (Kirsten et al., 2022) and FRB 20190520B (Niu et al., 2022). It has also discovered a non-repeating FRB 20190614D commensaly during a VLA observation (Law et al., 2020). In this paper, we report on the discovery and localization of a Galactic radio bursting source, associated with a Fermi -ray source, discovered by realfast.
@@
2 Data and Burst Detections
@@ 2.1 Discovery and Fermi Coincidence
The source in this paper, called J1818–1531 hereafter, was detected on 2019 September 2 (MJD 58728) as a part of VLA program SC1046, which imaged the regions near unidentified Fermi -ray sources in the inner galaxy. VLA was in the A configuration with 27 antennas and 351 baselines. The observation was done at VLA L-Band at a center frequency of 1.4 GHz. The total bandwidth of 1024 MHz was divided into 16 spectral windows, each with 64 channels having 1 MHz channel bandwidth. The main goal of the project was to identify steep-spectrum radio counterparts associated with the Fermi sources. realfast, a commensal fast-transient detection system operating on the VLA; (Law et al., 2015)), was operating during this project. While the telescope was pointed at RA=18:18:37.91 and DEC=-15:33:41.39 for one of the SC1046 targets, the realfast system detected five pulses at 1.4 GHz. The measured properties of these pulses are discussed below and presented in Table 1, and had an average DM of around 1016 pc cm. The pulse position was coincident with Fermi 95% -ray error ellipses of 3FGL J1818.7-1528 using 4 years of Fermi data (Acero et al., 2015), and 4FGL J1818.6-1533 using eight years of Fermi data (Abdollahi et al., 2020). The -ray source 4FGL J1818.6-1533 has two analysis flags (8196), making it part of the group of soft Galactic unassociated sources with higher flux uncertainty due to changes with older model or analysis. The DM-derived distance is 10.9 kpc using the NE2001 electron density model (Cordes & Lazio, 2002), and 5.8 kpc using the YMW16 model (Yao et al., 2017). This puts the source at a distance of 3.7 (3.03) kpc away from the Galactic Center for NE2001 (YMW16) distances.
2.1.1 Localization of Bursts
@@
The realfast search system was running alongside the SC1046 VLA observations. The visibilities are correlated commensally and sampled at a 10-ms resolution. These visibilites are then given to the graphics processing unit cluster on which the search pipeline rfpipe applies calibration, dedisperses, and forms images at many trial widths and DMs. The 8 fluence threshold limit of a 10-ms image is 0.29 Jy ms for the L band. If the image S/N is greater than the threshold, the fast-sampled visibilities, 2 to 5 s that include the candidate, are recorded. The frequency-time data averaged over visibilities for each candidate is processed and classified using the machine learning classifier Fetch (Agarwal et al., 2020). The realfast image of the candidates is convolved with the point spread function and is calibrated in real-time. The real-time images are made with several assumptions, including coarse DM grid, non-optimal image size, simpler calibration, etc. To resolve this, we used the raw visibilities dedispersed at the real-time detected DM to make burst images with CASA. The visibilities in the science data model (SDM) format were converted to measurement set (MS) format using the CASA tasks importasdm. The MS files were clipped for zeros using standard CASA flagging flagdata. We also applied Hanning-Smoothing and tfcrop to remove RFI from the data. We used the CASA calibration tables from the NRAO Archive for this observation and applied them to the raw data using the task applycal. The bright quasar 3C286 was used for flux and bandpass calibration. The phase calibrator J1911-2006 was observed for 90 seconds at regular intervals to calibrate complex gain fluctuations over time. The calibrated bursts were imaged using the CASA task tclean and the task imfit was used to fit an elliptical Gaussian to the burst in the image to get the centroid position, flux density, and the 1 image plane uncertainties.
@@
To determine the accuracy of the astrometric reference frame in our VLA observation (and ultimately to report the burst position), we created a deep image using the whole VLA pointing on this field, and ran PyBDSF111https://github.com/lofar-astron/PyBDSF to extract radio sources from it. This resulted in the detection of 99 radio sources. We then selected bright, compact radio sources using the following criteria: 1) The peak intensity per beam of the source (in Jy/beam) should be 0.7 times greater than the total integrated flux density of the source (in Jy) in 1.5 GHz images, 2) the S/N of the source (ratio of peak intensity and the root-mean-square of the background) should be greater than 5. There were 23 radio sources after the cut-off was applied. The positions of the radio sources had an average image-plane statistical error as reported by PyBDSF of 0.06 in RA and 0.11 in DEC.
@@
We then cross-matched the radio point sources with the optical PanSTARRS-DR2 catalog, which is referenced to the GAIA2-based astrometric reference frame. We identified radio/optical associations searching for the nearest PanSTARRS-DR2 source from each radio position, finding 23 optical counterparts in the PanSTARRS-DR2 catalog with a maximum separation from the radio component of 2.72.
@@
To determine whether there is a systematic offset between our radio imaging and the PanSTARRS-DR2 source catalog, we subtracted the coordinates of the radio sources from the matched coordinates of the optical counterparts. We then averaged the offset values to determine a systematic relative offset of and . The standard deviation in is 1.4 and in is 1.18.
@@
The average burst position at L-Band is RA and DEC. The average image-plane statistical error on the burst positions is and as reported by the CASA image-plane fitting function imfit. To represent the full positional error on the bursts, we add in quadrature sum these burst statistical errors with the Pan-STARRS offset, arriving at a final positional error of and . The realfast burst properties are reported in Table 1.
RA | DEC | DM | Flux | S/N | ||
(deg) | () | (deg) | () | () | (Jy) | |
274.6438328 | 0.02 | -15.5261568 | 0.06 | 1021.2 | 31.3 | |
274.6438538 | 0.02 | -15.5261619 | 0.04 | 1007.3 | 28.2 | |
274.6438212 | 0.03 | -15.5261553 | 0.05 | 1016.5 | 23.1 | |
274.6435122 | 0.04 | -15.5261382 | 0.12 | 1030.3 | 8.6 | |
274.6434000 | 0.04 | -15.5260776 | 0.10 | 1002.7 | 8.4 |
@@
realfast searches an extremely large number of samples and has a standard Gaussian noise threshold of 10. The last two bursts in table not used for positional analysis but are reported here because of their general proximity in sky position and DM.
@@ 2.2.1 VLA/realfast
To reduce the contributions from scattering and attempt to observe the intrinsic pulse structure, we carried out a higher-frequency observation, observing for three hours with the VLA observation at S-Band (2000-4000 MHz frequency) in phased-array mode under program code 19B-313. We configured the observation to observe in 64 sub-bands each of 32 MHz width, with 64 channels per subband, 100s time resolution, recording dual polarization. The VLA configuration was AD.
@@
We carried out a second set of GBT observations (AGBT 20B-407) on 6th August 2020 at 22:11:35.00 UTC for a total of 5 hours and 40 min on the source. This observation used the L-Band receiver with a center frequency of 1.5 GHz and a bandwidth of 800 MHz (4096 channels). We used the VEGAS pulsar backend in 8-bit format with a sampling time of 81s, and a channel frequency resolution of 195 kHz. We recorded the Full Stokes data in the IQUV format. A bright quasar, J1445+0958, used an injected signal for flux calibration, and a 1-minute noise diode scan was done on-source for polarization calibration. Hereafter we refer to this as “Epoch 2L.”
@@ 2.3 Single pulse searches
Here we describe the single pulse search done on the phased-array VLA S-band data and GBT data. The epoch 1 GBT data only had a usable bandwidth of 534 MHz, as the remaining parts of the band were either automatically filtered by the observing system or was manually flagged due to radio frequency interference (RFI). The GBT records data in Psrfits format, which we converted to Filterbank format using your_writer.py222https://github.com/thepetabyteproject/your. Custom RFI filters, which Savitzky–Golay (SG) and Spectral Kurtosis filter (Nita & Gary, 2010) with 4- threshold and an SG filter window of 15 MHz were used during the writing process so that the converted Filterbank was RFI cleaned. The VLA data was already recorded in Filterbank format. Single pulse search was done using the python code your_heimdall.py, which runs Heimdall (Barsdell, 2012) on the data, with a maximum boxcar width of 50 ms. The VLA data was dispersed at trial DMs between 990-1130 pc cm. The GBT epoch-1 data was searched in a wide range of DMs from 600-2500 pc cm. The epoch-2 1.5 GHz data was searched in a DM range of 600-2000 pc cm, and the 6 GHz data was searched in a DM range of 800-1200 pc cm. The Heimdall candidates were classified into real astrophysical signals both by the machine learning classifier Fetch (Agarwal et al., 2020) and by visual inspection of the candidate plots. A total of seven bursts were detected in the GBT data at 1.5 GHz, and three bursts were detected in the 6 GHz GBT data, above a signal-to-noise (S/N) ratio of 7. No bursts were detected in the phased VLA S-band data. Fig. 1 shows the bursts detected during the GBT observations.
@@
ID | MJD | S/N | DM | Fluence | Width | |
() | () | (ms) | (ms) | |||
L1 | 58930.418696580(8) | 24 | 1008.634.23 | 2.530.307 | 16.31.96 | 25.42.17 |
L2(a) | 58930.51873115(1) | 24 | 1021.463.87 | 3.780.667 | 16.42.87 | 26.82.87 |
L2(b) | 58930.51873150(2) | 24.26.54 | ||||
L3 | 58930.546647041(9) | 7 | 1002.293.62 | 0.7150.220 | 8.222.51 | 25.64.06 |
L4 | 58930.551574405(5) | 26 | 1011.862.71 | 4.190.392 | 13.51.27 | 29.41.54 |
L5(a) | 59087.980302006(6) | 1.060.15 | 8.661.91 | |||
L5(b) | 59087.980301320(2) | 197 | 1013.750.37 | 4.670.341 | 14.60.899 | 18.70.722 |
L5(c) | 59087.98030161(1) | 11.80.404 | 52.91.76 | |||
L6(a) | 59088.014787158(5) | 0.7460.32 | 3.251.8 | |||
L6(b) | 59088.01478710(1) | 141 | 1011.390.37 | 5.851.03 | 18.51.85 | 18.50.826 |
L6(c) | 59088.01478734(2) | 2.810.743 | 17.43.59 | |||
L7 | 59088.052215423(1) | 21 | 1014.542.28 | 1.320.0557 | 24.22.57 | 16.02.57 |
C1(a) | 59093.035770803(1) | 0.06720.00308 | 4.760.297 | |||
C1(b) | 59093.035770924(1) | 7 | 1003.68 4.23 | 0.2760.00592 | 6.530.194 | 0.9270.101 |
C1(c) | 59093.0357709811(7) | 0.04620.00537 | 1.640.204 | |||
C1(d) | 59093.0357710179(6) | 0.09860.00394 | 2.510.191 | |||
C2 | 59093.120417263(1) | 8 | 1002.784.24 | 0.09380.00911 | 10.7153.536 | 3.9012.1 |
C3 | 59093.155093193(3) | 11 | 999.934.24 | 0.01970.00159 | 2.240.456 | 0.8310.452 |

### v2

@@ Abstract
We report on the detection of radio bursts from the Galactic bulge using the real-time transient detection and localization system, realfast. The pulses were detected commensally on the Karl G. Jansky Very Large Array during a survey of unidentified Fermi -ray sources. The bursts were localized to subarcsecond precision using realfast fast-sampled imaging. Follow-up observations with the Green Bank Telescope detected additional bursts from the same source. The bursts do not exhibit periodicity in a search up to periods of 480 s, assuming a duty cycle of 20%. The pulses are nearly 100% linearly polarized, show circular polarization up to 12%, and exhibit variable scattering on timescales of months. The arcsecond-level realfast localization links the source confidently with the Fermi -ray source and places it nearby (though not coincident with) an XMM-Newton X-ray source. Based on the source’s overall properties, we discuss various options for the nature of this object and propose that it could be a young pulsar, a magnetar, or a binary pulsar system.
@@ 1 Introduction
Astrophysical transients are events whose duration ranges between milliseconds to years. They inform us about the most variable and energetic events in the Universe. Radio transients are typically described as “fast” if they last for 1 s (Law et al., 2018a). Some of the fast radio transients, including pulsars (Hewish et al., 1968), rotating radio transients (RRATs; McLaughlin et al. 2006), magnetar bursts, and fast radio bursts (FRBs) (Lorimer et al., 2007), are millisecond-duration emissions that are energetic enough to invoke coherent emission mechanisms to describe their high brightness temperature ( K) and luminosities. Most of these are produced by highly magnetized neutron stars, even though the origin of FRBs is highly debated. Pulsars, RRATs, and magnetars have a rotational period, while FRBs do not. FRBs are also many orders of magnitude brighter and more energetic than Galactic pulsars, magnetars, or RRATs, with the exception of FRB20200120E which has emitted bursts less luminous than FRB-like event FRB 200428 from the Galactic magnetar SGR 1935+2154. Nevertheless, these transients are unique probes in understanding the baryonic content of the Milky Way interstellar medium (ISM) and the intergalactic medium (IGM), in the case of FRBs (Macquart et al., 2020), along their line-of-sight (LOS). Regular timing of millisecond pulsars (MSPs) has been used to detect stochastic gravitational wave background in the Universe (Hellings & Downs, 1983; Agazie et al., 2023). Multiple ways have been proposed to detect these radio transients, including all-sky blind searches. One such method is to look for radio counterparts in the known high-energy sources, as some of these sources, like pulsars and magnetars, are established high-energy emitters.
@@
Fermi is a space mission that studies the cosmos in the energy range 10 keV–300 GeV. The imaging telescope on Fermi called the Large Area Telescope (LAT) has a wide field of view (8000 cm-2) and an angular resolution of . Fermi-LAT (4FGL-DR4) reported about 7195 point sources in their latest data release (Ballet et al., 2023). About 2427 (one-third) of these Fermi sources don’t have counterparts in any other electromagnetic regime. The origin and nature of these unassociated sources remain a mystery. A large fraction of these are eventually expected to be associated with the largest associated source classes, active galactic nuclei or pulsars. Recently, there appears a third class of so-called soft Galactic unassociated sources that are found in high-density regions of the Galactic plane. For a more detailed discussion of the unassociated gamma-ray source population see (Abdollahi et al. (2022)). Fermi-LAT has detected at least 297 -ray pulsars, 70% of which are radio-loud, and MSPs and young pulsars dominate this sample (Smith et al., 2023). Therefore, searching for new pulsars in the Fermi sources remains a reasonable case.
@@
The potential for precise localization capabilities make radio interferometers well-suited for searching for radio transients. Radio imaging-based pulsar searches have led to the discovery of the first isolated MSPs (Erickson, 1980) and the first globular cluster pulsar (Hamilton et al., 1985). This technique is based on the assumption that pulsars are compact and steep spectrum radio sources. Such searches have successfully led to the discovery of many MSPs and normal pulsars (Frail et al., 2016; Bhakta et al., 2017).
@@
realfast is a real-time commensal transient search system deployed at the Karl G. Jansky Very Large Array (JVLA) (Law et al., 2015, 2018b). It makes real-time interferometric images on visibility data sampled at 10 ms duration. This has the advantage of simultaneously detecting and localizing the transient to sub-arcsecond precision. realfast has so far played a crucial role in the localization and host galaxy determination of many FRBs, including the first repeating FRB 20121102A (Chatterjee et al., 2017), FRB 20180916B (Aggarwal et al., 2020), FRB 20180301A (Bhandari et al., 2022), FRB 20201124A (Ravi et al., 2022), FRB 20200120E (Kirsten et al., 2022) and FRB 20190520B (Niu et al., 2022). It has also discovered a non-repeating FRB 20190614D in commensal during a VLA observation (Law et al., 2020). In this paper, we report on the discovery and localization of a Galactic radio bursting source, associated with a Fermi -ray source, discovered by realfast.
@@
2 Data and Burst Detection
@@ 2.1 Discovery and Fermi Coincidence
The source in this paper, called J1818–1531 hereafter, was detected on 2019 September 2 (MJD 58728) as a part of VLA program SC1046, which imaged the regions near unidentified Fermi -ray sources in the inner Galaxy. The VLA was in the A configuration with 27 antennas and 351 baselines. The observation was done at VLA L-Band at a center frequency of 1.4 GHz. The total bandwidth of 1024 MHz was divided into 16 spectral windows, each with 64 channels having 1 MHz channel bandwidth. The main goal of the project was to identify steep-spectrum radio counterparts associated with the Fermi sources. While the telescope was pointed at RA=18:18:37.91 and DEC=-15:33:41.39 for one of the SC1046 targets, the realfast system detected five pulses at 1.4 GHz. The measured properties of these pulses are discussed below and presented in Table 1, and had an average dispersion measure (DM) of around 1016 pc cm-3. The pulse position was coincident with Fermi 95% -ray error ellipses of 3FGL J1818.7-1528 using 4 years of Fermi data (Acero et al., 2015), and 4FGL J1818.6-1533 using eight years of Fermi data (Abdollahi et al., 2020). The -ray source 4FGL J1818.6-1533 has two analysis flags (8196), making it part of the group of soft Galactic unassociated sources with higher flux uncertainty due to changes with older model or analysis. The DM-derived distance is 10.9 kpc using the NE2001 electron density model (Cordes & Lazio, 2002), and 5.8 kpc using the YMW16 model (Yao et al., 2017). This puts the source at a distance of 3.7 (3.03) kpc away from the Galactic Center for NE2001 (YMW16) distances.
2.1.1 Burst searching and localization
@@
The realfast search system was running alongside the SC1046 VLA observations. The visibilities are correlated commensally and sampled at a 10-ms resolution. These visibilities are then given to the graphics processing unit cluster on which the search pipeline rfpipe applies calibration, dedisperses, and forms images at many trial widths and DMs. The 8 fluence threshold limit of a 10-ms image is 0.29 Jy ms for the L-band. If the image S/N is greater than the threshold, the fast-sampled visibilities, 2 to 5 s that include the candidate, are recorded. The frequency-time data averaged over visibilities for each candidate is processed and classified using the machine learning classifier Fetch (Agarwal et al., 2020). The realfast image of the candidates is convolved with the point spread function and is calibrated in real-time. The real-time images are made with several assumptions, including coarse DM grid, non-optimal image size, simpler calibration, etc. To resolve this, we used the raw visibilities dedispersed at the real-time detected DM to make burst images with CASA. The visibilities in the science data model (SDM) format were converted to measurement set (MS) format using the CASA tasks importasdm. The MS files were clipped for zeros using standard CASA flagging flagdata. We also applied Hanning-Smoothing and tfcrop to remove RFI from the data. We used the CASA calibration tables from the NRAO Archive for this observation and applied them to the raw data using the task applycal. The bright quasar 3C286 was used for flux and bandpass calibration. The phase calibrator J1911-2006 was observed for 90 seconds at regular intervals to calibrate complex gain fluctuations over time. The calibrated bursts were imaged using the CASA task tclean and the task imfit was used to fit an elliptical Gaussian to the burst in the image to get the centroid position, flux density, and the 1 image plane uncertainties.
@@
To determine the accuracy of the astrometric reference frame in our VLA observation (and ultimately to report the burst position), we created a deep image using the whole VLA pointing on this field, and ran PyBDSF111https://github.com/lofar-astron/PyBDSF to extract radio sources from it. This resulted in the detection of 99 radio sources. We then selected bright, compact radio sources using the following criteria: 1) The peak intensity per beam of the source (in Jy/beam) should be 0.7 times greater than the total integrated flux density of the source (in Jy) in 1.5 GHz images, 2) the S/N of the source (ratio of peak intensity and the root-mean-square of the background) should be greater than 5. There were 23 radio sources after the cut-off was applied. The positions of the radio sources had an average image-plane statistical error as reported by PyBDSF of 0.06′′ in RA and 0.11′′ in DEC.
@@
We then cross-matched the radio point sources with the optical PAN-STARRS DR2 catalog, which is referenced to the GAIA2-based astrometric reference frame. We identified radio/optical associations searching for the nearest PAN-STARRS-DR2 source from each radio position, finding 23 optical counterparts in the PAN-STARRS-DR2 catalog with a maximum separation from the radio component of 2.72′′.
@@
To determine whether there is a systematic offset between our radio imaging and the PAN-STARRS-DR2 source catalog, we subtracted the coordinates of the radio sources from the matched coordinates of the optical counterparts. We then averaged the offset values to determine a systematic relative offset of and . The standard deviation in is 1.4′′ and in is 1.18′′.
@@
The average burst position at L-Band is RA and DEC. The average image-plane statistical error on the burst positions is and as reported by the CASA image-plane fitting function imfit. To represent the full positional error on the bursts, we add in quadrature sum these burst statistical errors with the PAN-STARRS offset, arriving at a final positional error of and . The realfast burst properties are reported in Table 1 and the dirty and clean images of the visibility images of the bursts are given in Figure. 1
ID | RA | DEC | DM | Flux | S/N | ||
(deg) | (′′) | (deg) | (′′) | () | (Jy) | ||
R1 | 274.64383 | 0.02 | -15.52615 | 0.06 | 1021.2 | 31.3 | |
R2 | 274.64385 | 0.02 | -15.52616 | 0.04 | 1007.3 | 28.2 | |
R3 | 274.64382 | 0.03 | -15.52615 | 0.05 | 1016.5 | 23.1 | |
R4 | 274.64351 | 0.04 | -15.52613 | 0.12 | 1030.3 | 8.6 | |
R5 | 274.64340 | 0.04 | -15.52607 | 0.10 | 1002.7 | 8.4 |
@@
realfast searches an extremely large number of samples and has a standard Gaussian noise threshold of 7. The last two bursts in the table not used for positional analysis but are reported here because of their general proximity in sky position and DM.
@@ 2.2.1 VLA/realfast
To reduce the contributions from scattering and attempt to observe the intrinsic pulse structure, we carried out a higher-frequency observation, observing for three hours with the VLA observation at S-Band (2000-4000 MHz frequency) in phased-array mode under program code 19B-313. We configured the observation to observe in 64 sub-bands each of 32 MHz width, with 64 channels per sub-band, 100s time resolution, recording dual polarization. The observation was during the VLA reconfiguration from A to D (AD)
@@
We carried out a second set of GBT observations (AGBT 20B-407) on 6th August 2020 at 22:11:35.00 UTC for a total of 5 hours and 40 min on the source. This observation used the L-Band receiver with a center frequency of 1.5 GHz and a bandwidth of 800 MHz (4096 channels). We used the VEGAS pulsar backend in 8-bit format with a sampling time of 81s, and a frequency resolution of 195 kHz. We recorded the Full Stokes data in the IQUV format. A bright quasar, J1445+0958 was observed for flux calibration, and a 1-minute noise diode scan was done on-source for polarization calibration. Hereafter we refer to this as “Epoch 2L.”
@@ 2.3 Single pulse searches
Here we describe the single pulse search done on the phased-array VLA S-band data and GBT data. The epoch 1 GBT data only had a usable bandwidth of 534 MHz, as the remaining parts of the band were either automatically filtered by the observing system or was manually flagged due to radio frequency interference (RFI). The GBT records data in Psrfits format, which we converted to Filterbank format using your_writer.py222https://github.com/thepetabyteproject/your. Custom RFI filters, with Savitzky–Golay (SG) and Spectral Kurtosis filter (Nita & Gary, 2010) with 4- threshold and an SG filter window of 15 MHz were used during the writing process so that the converted Filterbank was RFI cleaned. The VLA data was already recorded in Filterbank format. Single pulse search was done using the python code your_heimdall.py, which runs Heimdall (Barsdell, 2012) on the data, with a maximum boxcar width of 50 ms and a detection threshold of 6. The VLA data was dispersed at trial DMs between 990-1130 pc cm-3. The GBT epoch-1 data was searched in a wide range of DMs from 600-2500 pc cm-3. The Epoch 2L data was searched in a DM range of 600-2000 pc cm-3, and the 6 GHz data was searched in a DM range of 800-1200 pc cm-3. The Heimdall candidates were classified into real astrophysical signals both by the machine learning classifier Fetch (Agarwal et al., 2020) and by visual inspection of the candidate plots. A total of seven bursts were detected in the GBT data at 1.5 GHz, and three bursts were detected in the 6 GHz GBT data, above a signal-to-noise (S/N) ratio of 7. No bursts were detected in the phased VLA S-band data. Fig. 2 shows the bursts detected during the GBT observations.
@@
is the reduced chi-squared value obtained for profile fitting.
ID | MJD | S/N | DM | Fluence | Width | ||
() | () | (ms) | (ms) | ||||
L1 | 58930.418696580(8) | 7 | 1008.634.23 | 0.2830.0528 | 16.31.96 | 25.42.17 | 1.05 |
L2(a) | 58930.51873115(1) | 11 | 1021.463.87 | 0.7010.139 | 16.42.87 | 26.82.87 | 1.35 |
L2(b) | 58930.51873150(2) | 24.26.54 | |||||
L3 | 58930.546648171(9) | 6 | 1002.293.62 | 0.1720.0598 | 8.222.51 | 25.64.06 | 0.838 |
L4 | 58930.551574405(5) | 18 | 1011.862.71 | 0.6620.0718 | 13.51.27 | 29.41.54 | 1.57 |
L5(a) | 59087.980302006(6) | 2.300.327 | 8.661.91 | ||||
L5(b) | 59087.980301320(2) | 33 | 1013.750.37 | 25.70.879 | 52.81.76 | 18.70.722 | 1.30 |
L5(c) | 59087.98030161(1) | 10.20.742 | 14.60.899 | ||||
L6(a) | 59088.014787158(5) | 1.650.726 | 3.261.81 | ||||
L6(b) | 59088.014787096(1) | 28 | 1011.390.37 | 12.52.28 | 18.11.79 | 18.50.826 | 1.08 |
L6(c) | 59088.01478733 (2) | 6.341.64 | 17.83.62 | ||||
L7 | 59088.052215423(1) | 7 | 1014.542.28 | 2.870.121 | 24.22.57 | 16.02.57 | 1.00 |
C1(a) | 59093.035770803(1) | 0.08160.00377 | 4.760.305 | ||||
C1(b) | 59093.035770924(1) | 66 | 1003.68 4.23 | 0.3340.00718 | 6.530.206 | 0.9270.101 | 2.33 |
C1(c) | 59093.0357709811(7) | 0.05620.00735 | 1.640.204 | ||||
C1(d) | 59093.0357710179(6) | 0.1190.0058 | 2.510.266 | ||||
C2 | 59093.120416971(2) | 9 | 1002.784.24 | 0.02960.00201 | 2.760.0958 | 0.5730.0105 | 0.947 |
C3 | 59093.155093193(3) | 13 | 999.934.24 | 0.02730.00221 | 2.240.456 | 0.8310.452 | 1.23 |